OpenStudy (anonymous):

there are about 10 ^11 stars in the milky way galaxy. if each star has about 10 planets orbiting around it,how many planets are there in the milky way galaxy?

OpenStudy (mrnood):

there are 10 times as many planets as stars so 10^11 * 10 Find a neater way to write that....

OpenStudy (anonymous):

Playing with units! 10^11 stars/galaxy 10 planets/star so: (10 planets/star)*(10^11 stars/galaxy) = 10*10^11 (planets/star)*(star/galaxy) cancel common units: = 10*10^11 (planets/galaxy)*(star/star) and multiplying constants: = 10^12 (planets/galaxy)

OpenStudy (anonymous):

so the answer is 10^12 ? just adding 1 more 0

OpenStudy (mrnood):

@felis From code of conduct Don't devalue the question/answer process! •Don't provide someone with just the answer - explain the process, and help guide them through understanding the problem. •Don't just provide the answer to a problem when someone else is in the middle of helping! But if you want to help, by all means, join in!

jaynator495 (jaynator495):

\huge\cal\color{lime}{1,000,000,000,000 LOL} the way you get it is 10^11 which is 10x10x10x10x10x10x10x10x10x10x10 then because theres another 10x thrown in there multiply the answer by 10

OpenStudy (mrnood):

whenever you multiply by 10 you add one more zero However this is more about exponents 10^1 * 10^11 = 10 ^(11+1) = 10 ^12

jaynator495 (jaynator495):

dang it didnt work... still helped him out lol

OpenStudy (anonymous):

@MrNood I wanted to show how unit cancellation can help inform thinking on word problems, was meaning to add to the understanding of the interpretation of the question :3

OpenStudy (anonymous):

ok thanks i got it! and felis the cancellation did help

OpenStudy (anonymous):

I'm glad! Remember, if you are ever confused about how to solve a problem, especially a word problem, try writing out the units for all the important numbers, and seeing how they fit together, it can really change how you see the problem!